OpenStudy (anonymous):

which is the equation for the line perpendicular to y= -5/3x + 11 1/3 ad containing p(-2, 3)

OpenStudy (aum):

What is the slope of the line y = -5/3x + 11 1/3 ?

OpenStudy (aum):

y = mx + b where 'm' is the slope. y = -5/3x + 11 1/3 So what is m?